OpenStudy (theedwardsfamily):

Jeanie drank 1/3 cup of juice and Harry drank 1/4 cup of juice. How much juice did they drink altogether?

OpenStudy (anonymous):

Add the fractions, first we need a common denominator, which is 12. \(\dfrac{1}{3} \rightarrow \dfrac{1 \times 4}{3 \times 4} \rightarrow \dfrac{4}{12}\) \(\dfrac{1}{4} \rightarrow \dfrac{1 \times 3}{4 \times 3} \rightarrow \dfrac{3}{12}\) Add them: \(\dfrac{4}{12} + \dfrac{3}{12} \rightarrow \dfrac{4 + 3}{12}\)

OpenStudy (anonymous):

Add 4 + 3 @TheEdwardsFamily

OpenStudy (legends):